two years ago the age of a man was four times the age of his son.3 years from now the age of the man will be 3 times the age of his son.find the age of each of them now

Answer by macston(5194)

two years ago the age of a man was four times the age of his son.3 years from now the age of the man will be 3 times the age of his son.find the age of each of them now
F=fathers age; S=son's age
F-2=4(S-2)
F-2=4S-8 Add 2 to each side
F=4s-6
F+3=3(S+3) Substitute for F
4S-6+3=3S+9
4S-3=3S+9 subtract 3S from each side,
S-3=9 Add 3 to each side
S=12 ANSWER 1: The son is 12 years old.
F=4S-6=4(12)-6=48-6=42 ANSWER 2: The father is 42 years old.
CHECK
F-2=4(S-2)
42-2=4(12-2)
40=4(10)
40=40